# Byam: Fixing Breaking Dependency Updates with Large Language Models

### v1

@@ 4.4 Language Models for Experimentation
We select the models we experiment of based on the results of RepairBench (Silva and Monperrus, 2024) and LiveCodeBench (Jain et al., 2024), which are established evaluation frameworks for LLMs in program repair. Based on these frameworks, we select the following five LLMs to use in our evaluation:
Table 3 shows a comparison of the five models according to different aspects like their provider, number of parameters, and input and output token limits.
@@ 5 Experimental Results
The highest success rate in this evaluation is achieved by o3-mini using $P_{8}$ (CoT + Buggy Line + API Diff), with 27%, meaning 28 out of the 103  original failing builds were completely repaired, with compilation and test execution suceeding. Overall, o3-mini had the highest build success rate across most prompts. Relatively high performance is also achieved by Deepseek V3 performing second-best in cases $P_{1}$, $P_{5}$, $P_{6}$ and Gemini-2.0-flash performing second-best in $P_{2}$, $P_{3}$, $P_{4}$, $P_{8}$
@@
Given that o3-mini has the highest build success rate, we further analyze the its fixes in Figure 11. The figure illustrates the number of breaking commits successfully fixed by o3-mini across prompts, and the overlap between prompts. It shows each prompt ($P_{1}$ through $P_{8}$) as a horizontally colored row. The right-side number indicates how many commits each prompt successfully fixes in total. Along the upper part of the diagram, the numbers indicate how many prompts overlap in fixing a particular set of breaking updates, and the numbers at the bottom of the diagram are the actual count of these commits. The width of each ’chunk’ represents how many breaking updates fall into those intersections between prompts. For example, the chunk labeled ‘8’ at the top and ‘16’ at the bottom indicates that there are 16 breaking updates successfully fixed by all eight prompts. Overall, $P_{8}$ fixed the most commits (28), while $P_{4}$ fixed the least breaking updates (20), confirming the importance of API Diff and Cot.
@@
Answer to RQ1: It is possible to fix real-world breaking updates with LLMs. o3-mini with prompt $P_{8}$ (CoT + Buggy Line + APIDiff) achieves the highest build success rate, fixing 27 out of the 103 original failing builds (28%). The addition of API Diff and Chains-of-thought makes a difference, yet a relatively small one. Open-source model Qwen2.5-32b-instruct, which is the smallest considered model, has the lowest success rate, as expected. Overall, the inherent capability of the model is more important than the prompting itself for fixing breaking updates.
@@
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@
It is clearly that Byam gets lots of partial fixes, in the double digits range. For example, when it fails to completely repair the build, o3-mini with $P_{4}$ (Buggy Line + API Diff) is able to fix 97 of the 239 original files with compilation failures (41%). All models are able to fix dozens of files.
At the prompt level, adding buggy lines and APIDiff information improves partial repair, confirming RQ1. For instance, Gemini-2.0-flash improves from 19% (48/251) with the baseline prompt ($P_{1}$) to 35% (87/251) with $P_{4}$, which shows that the addition of APIDiff and buggy lines in the configurations improves the performance of the model.
This is more evidence that the inclusion of APIDiff in the prompts assists the model to understand the nature of the breakage more clearly and to generate a more appropriate fix. On the other hand, GPT-4o-mini shows low values with the CoT prompt, achieving only 14% for $P_{5}$. However, its performance improves notably when using APIDiff and buggy line information. This highlights how reasoning plays a significant role in influencing model performance. However, the addition of CoT improves performance in Qwen2.5-32b-instruct, from 16% in $P_{1}$ to 24% in $P_{5}$ and $P_{6}$. This indicates that CoT positively impacts the performance for fixing error files in specific models.
@@
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@
We summarize the results in Table 7, where we show the median value of the REF. Across all breaking builds, o3-mini performs the best, delivering the highest median REF on every prompt, peaking at 93.33% with $P_{4}$ (Buggy Line and APIDiff). It also shows promising stability, never dropping below 66.67 %, which shows that the model is robust enough that providing it with basic information ($P_{1}$-$P_{5}$) or more advanced reasoning and contextual information ($P6$-$P8$ ) will not result in new errors.
@@
When we consider the results from the prompts viewpoint, we see that the improvement of $P_{2}$ (Buggy Line) is consistently better, boasting the results of four of the five models. On the contrary, $P_{5}$ (CoT) is the least consistently effective, helping only o3-mini. This indicates that adding the exact buggy line is a valuable prompt addition to most LLMs. CoT has an added value only when the model supports reasoning, such as o3-mini.
@@
Prompt ID Deepseek V3 Gemini 2.0-flash Gpt 4o-mini o3 mini Qwen2.5 32b-instruct
@@
Answer to RQ3: Our results show that o3-mini is the model that introduces the least new errors. Prompt-wise, providing the buggy line ($P_{2}$) yields the most gains across all LLMs to avoid new error introduction. Models with reasoning abilities are the ones which introduce the least new errors, benefiting more from the added contextual information.
@@ 7 Related Work
Breaking dependency updates is a common problem in software development. (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020) investigate the reasons behind the developer’s decision to introduce updates that lead to breaking dependency updates. These reasons include introducing new features, refactoring the code, fixing bugs, and addressing security vulnerabilities. Hejderup et al. analyze how changes in method behavior trigger incompatibilities despite preserving the API contract. Such changes are difficult to detect by static analysis and often only become apparent at runtime (Hejderup and Gousios, 2021). (Venturini et al., 2023) emphasized the challenges developers face in updating client code after dependency updates. Researchers investigated various approaches to address this issue, where some investigated detecting breaking dependency updates (Brito et al., 2018a ; Mujahid et al., 2020). (Reyes et al., 2024a ) recently introduced an automated approach to explain the breaking in the client code after the version update. Our approach differs from the previous studies as it uniquely addresses compilation failures caused by breaking updates through fixes generated using LLM. Byam originally exploits contextual information such as APIDiff and buggy lines to augment prompts, a dimension unexplored in previous studies.
@@
Our approach differs from previous research as it focuses on fixing Java compilation errors across build, file, and error levels. The most closely related work to our study is presented in the recently published work by (Frunkte and Krinke, 2025). The authors introduce two new approaches to fix breaking dependency updates using LLMs: a zero-shot prompting method and an agentic system. Our approach differs in the depth of evaluation and prompt design. We propose a detailed evaluation at the file and error level, measuring success rate at both levels. Our work explores a wider set of prompt variants, analyzes the impact of prompt design choices, including chain reasoning, which is not addressed by (Frunkte and Krinke, 2025).
@@ Acknowledgements.
This work was supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), the WebInspector project funded by Swedish Research Council (VR), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.

### v2

@@ 4.4 Language Models for Experimentation
We select the models we experiment with based on the results of RepairBench (Silva and Monperrus, 2024) and LiveCodeBench (Jain et al., 2024), which are established evaluation frameworks for LLMs in program repair. Based on these frameworks, we select the following five LLMs to use in our evaluation:
Table 3 shows a comparison of the five models according to different aspects like their provider, number of parameters, and input and output token limits. Our study deliberately focuses on small and medium-sized, cost-effective models to enable large-scale, reproducible experimentation across many real breaking dependency updates and projects. To reduce sampling noise, we set the temperature to 0, which minimizes randomness; the remaining limitations of this choice are discussed in Subsection 7.1. This design choice aims to balance scalability, cost, and replicability in our evaluation.
@@ 5 Experimental Results
The highest success rate in this evaluation is achieved by o3-mini using $P_{8}$ (CoT + Erroneous Line + API Diff), with 27%, meaning 28 out of the 103  original failing builds were completely repaired, with compilation and test execution succeeding. Overall, o3-mini had the highest build success rate across most prompts. Relatively high performance is also achieved by Deepseek V3 performing second-best in cases $P_{1}$, $P_{5}$, $P_{6}$ and Gemini-2.0-flash performing second-best in $P_{2}$, $P_{3}$, $P_{4}$, $P_{8}$
@@
Given that o3-mini has the highest build success rate, we further analyze its fixes in Figure 11. The figure illustrates the number of breaking commits successfully fixed by o3-mini across prompts, and the overlap between prompts. It shows each prompt ($P_{1}$ through $P_{8}$) as a horizontally colored row. The right-side number indicates how many commits each prompt successfully fixes in total. Along the upper part of the diagram, the numbers indicate how many prompts overlap in fixing a particular set of breaking updates, and the numbers at the bottom of the diagram are the actual count of these commits. The width of each ’chunk’ represents how many breaking updates fall into those intersections between prompts. For example, the chunk labeled ‘8’ at the top and ‘16’ at the bottom indicates that there are 16 breaking updates successfully fixed by all eight prompts. Overall, $P_{8}$ fixed the most commits (28), while $P_{4}$ fixed the least breaking updates (20), confirming the importance of API Diff and Cot.
@@
Answer to RQ1: It is possible to fix real-world breaking updates with LLMs. o3-mini with prompt $P_{8}$ (CoT + Erroneous Line + APIDiff) achieves the highest build success rate, fixing 27 out of the 103 original failing builds (28%). The addition of API Diff and Chains-of-thought makes a difference, yet a relatively small one. Open-source model Qwen2.5-32b-instruct, has the lowest success rate, as expected. Overall, the inherent capability of the model is more important than the prompting itself for fixing breaking updates.
@@
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@
It is clearly that Byam gets lots of partial fixes, in the double digits range. For example, when it fails to completely repair the build, o3-mini with $P_{4}$ (Erroneous Line + API Diff) is able to fix 97 of the 239 original files with compilation failures (41%). All models are able to fix dozens of files.
At the prompt level, adding erroneous lines and APIDiff information improves partial repair, confirming RQ1. For instance, Gemini-2.0-flash improves from 19% (48/251) with the baseline prompt ($P_{1}$) to 35% (87/251) with $P_{4}$, which shows that the addition of APIDiff and erroneous lines in the configurations improves the performance of the model.
This is more evidence that the inclusion of APIDiff in the prompts assists the model to understand the nature of the breakage more clearly and to generate a more appropriate fix. On the other hand, GPT-4o-mini shows low values with the CoT prompt, achieving only 14% for $P_{5}$. However, its performance improves notably when using APIDiff and erroneous line information. This highlights how reasoning plays a significant role in influencing model performance. However, the addition of CoT improves performance in Qwen2.5-32b-instruct, from 16% in $P_{1}$ to 24% in $P_{5}$ and $P_{6}$. This indicates that CoT positively impacts the performance for fixing error files in specific models.
@@
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@
We summarize the results in Table 7, where we show the median value of the REF. Across all breaking builds, o3-mini performs the best, delivering the highest median REF on every prompt, peaking at 93.33% with $P_{4}$ (Erroneous Line and APIDiff). It also shows promising stability, never dropping below 66.67 %, which shows that the model is robust enough that providing it with basic information ($P_{1}$-$P_{5}$) or more advanced reasoning and contextual information ($P6$-$P8$ ) will not result in new errors.
@@
When we consider the results from the prompts viewpoint, we see that the improvement of $P_{2}$ (Erroneous Line) is consistently better, boasting the results of four of the five models. On the contrary, $P_{5}$ (CoT) is the least consistently effective, helping only o3-mini. This indicates that adding the exact erroneous line is a valuable prompt addition to most LLMs. CoT has an added value only when the model supports reasoning, such as o3-mini.
@@
Prompt
ID
Deepseek
V3
Gemini
2.0-flash
Gpt
4o-mini
o3
mini
Qwen2.5
32b-instruct
@@
Answer to RQ3: Our results show that o3-mini is the model that introduces the least new errors. Prompt-wise, providing the erroneous line ($P_{2}$) yields the most gains across all LLMs to avoid new error introduction. Models with reasoning abilities are the ones which introduce the least new errors, benefiting more from the added contextual information.
6 Discussion
In this section, we reflect on our findings and their implications for both future research and practical use. Subsection 6.1 compares Byam with prior work, situating our results in relation to existing approaches. Subsection 6.2 analyzes the main causes of unsuccessful repairs , highlighting the limits of fixes. Finally, Subsection 6.3 discusses how Byam can be integrated into developer workflows and outlines promising directions for future wrk.
6.1 Comparison with Prior Work
In Frunkte and Krinke, (2025), the authors introduce two new approaches to fix breaking dependency updates using LLMs: a zero-shot prompting method and an agent system. Their agent iteratively attempts up to 30 fixes per case. The authors process the Bump dataset to produce two distinct evaluation data sets, the first set is considered as “light slice” with 65 projects with breaking dependency updates caused by only one file, and a “full slice” of 140 projects that has multiple files causing breaking in build. The authors test the zero-shot approach on the light-slice dataset while they test the agent system on both datasets. The zero-shot baseline uses no prompt engineering or design and is evaluated only on the light-slice dataset. The agent system, by contrast, is evaluated on both slices and operates iteratively, making up to 30 repair attempts per case. Importantly, the study does not report single-iteration success rates for the agent system, as its evaluation focuses on cumulative results across multiple trials.
Byam’s key advantage over the work in Frunkte and Krinke, (2025) is its context-rich prompting, evident in being able to achieve better performance using a less complicated approach . Byam fixed full builds that can consist of multiple file edits, achieving a success rate of 27% using the richest prompt (P8). Even when the full build does not succeed, Byam achieves a 78% fix rate of the individual compilation errors.
In contrast, the agent approach in Frunkte and Krinke, (2025) fixes a maximum of 23% of builds in the full slice and a maximum of 19% in the light-slice, with the zero-shot approach achieving a maximum success rate of 19%. We note that Frunkte and Krinke, (2025) does not report the average trials for fixes but allow iteration for up to 30 trials.
To provide a direct comparison of identical breaking dependency updates, we analyze the intersection of breaking dependency updates evaluated by both studies. The intersection between the 140 (“full slice”) breaking updates presented by Frunkte and Krinke, (2025) and the 103 breaking updates analyzed here yields 97 common breaking dependency updates. When evaluating Byam with o3-mini using the P8 configuration on these 97 shared cases, Byam achieves a 27% (26/97) success rate, compared to the agent approach’s 20% (19/97) performance on the same breaking dependency updates. This controlled comparison on the same failing builds eliminates dataset variability and demonstrates that our contextualized prompting strategy outperforms the agent-based approach.
To sum up, the key strength and novelty of Byam is contextualize the breaking update problem in the prompt with essential information: by considering the line that is causing the errors, structuring the API diffs to the parts relevant to the project under investigation, and utilizing chain-of-thought prompting.
6.2 Unsuccessful Repairs and Test Failures
Our work aims at repairing compilation failures. Even when full builds are not repaired, Byam often generates partial repairs that provide practical assistance. For example, with the o3-mini model, Byam fixes 78% of individual errors and 41% of erroneous files in otherwise failing builds. This suggests that many builds fail only due to a small number of remaining issues, and Byam can significantly reduce developer effort by handling the bulk of the errors automatically.
Yet, once compilation failures are resolved, we have sometimes test failures, indicating that semantic changes have occurred in the library. For example, in project pay-adminusers, the dependency update of logback-classic from version 1.2.11 to 1.4.5 fails due to compilation error cannotaccessorg.slf4j.spi.LoggingEventAwareclassfilefororg.slf4j.spi.LoggingEventAwarenotfound. The compilation error is fixed by replacing the class ch.qos.logback.classic.Logger with the generic slf4j interface org.slf4j.Logger. However, during testing, the method doAppend() was never invoked, resulting in the error Wantedbutnotinvoked:mockLogAppender.doAppend(). This is an example where the updates by the LLM fix the compilation failure, but result in a test failure.
There are two ways in which we can interpret a test failure outcome after automated fixing by LLMs: (1) the LLM-generated code has introduced logical errors that pass compilation but fail functional tests, or (2) some unmodified parts of the code interact with the updated dependency, which has changed semantics. Distinguishing between these scenarios is a challenging and important direction for future work.
6.3 Workflow Integration and Future Work
Applications in the field of tools like Byam require integration. We note that there exist Pull Request (PR) bots, which create pull requests to update dependencies (Dependabot, Renovate). Byam could be integrated into such systems to further automate the process of dependency update. For example, if a PR for a dependency update causes a break in the Continuous Integration (CI) environment, an integrated tool could attempt to fix the breaking dependency update. Such an end-to-end solution to dependency updates would further reduce developers’ effort in keeping their dependencies up-to-date, thus making their projects more secure and reliable.
Future work should explore several directions. First, addressing semantic failures may require incorporating additional context from test suites or integrating program analysis to reason beyond syntax. Second, as more powerful reasoning-oriented LLMs emerge, they may further improve repair rates, particularly for complex or multi-file repairs. Finally, while our evaluation focused on Java/Maven projects, the core idea of contextual prompting is generalizable and could be applied to other ecosystems.
@@ 8 Related Work
Breaking dependency updates is a common problem in software development. (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020) investigate the reasons behind the developer’s decision to introduce updates that lead to breaking dependency updates. These reasons include introducing new features, refactoring the code, fixing bugs, and addressing security vulnerabilities. Hejderup et al. analyze how changes in method behavior trigger incompatibilities despite preserving the API contract. Such changes are difficult to detect by static analysis and often only become apparent at runtime (Hejderup and Gousios, 2021). (Venturini et al., 2023) emphasized the challenges developers face in updating client code after dependency updates. Researchers investigated various approaches to address this issue, where some investigated detecting breaking dependency updates (Brito et al., 2018a ; Mujahid et al., 2020). (Reyes et al., 2024a ) recently introduced an automated approach to explain the breaking in the client code after the version update. Our approach differs from the previous studies as it uniquely addresses compilation failures caused by breaking updates through fixes generated using LLM. Byam originally exploits contextual information such as APIDiff and erroneous lines to augment prompts, a dimension unexplored in previous studies.
@@
Our approach differs from previous research as it focuses on fixing Java compilation errors across build, file, and error levels.
The most closely related work to our study is presented in parallel research by  (Frunkte and Krinke, 2025). The authors explore automated build repair using either zero-shot prompting or agent-based iterative repair loops without explicit prompt engineering, achieving at most 23% repair on the full-slice dataset. Byam improves on this by incorporating contextual information, reaching 27% success on the same dataset of builds in a single iteration and fixing 78% of individual compilation errors.
@@ Acknowledgements.
This work was supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), the WebInspector project funded by Swedish Research Council (VR), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Some computation was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS).